You have 2.2 mol Xe and 4.0 mol F₂, but when you carry out the reaction you end up with only 0.25 mol XeF₄. What is the percent yield of this experiment?

Xe(g) + 2 F₂ (g) → XeF₄ (g)


1
Expert's answer
2022-05-31T07:43:03-0400

Xe + 2 F2 = XeF4

According to the above reaction, 1 mole of Xe reacts with 2 moles of F2 to form 1 mole of XeF4

2mol F2 => 1mol XeF4

4mol F2 => X

X = 4 * 1 / 2 = 2 mol XeF4

if the reaction proceeds with 100% efficiency, 2 mol XeF4 is formed. We find the yield of 0.25 mol of XeF4

2mol — 100%

0,25mol — X

X = 0,25 * 100 / 2 = 12,5 %

ANSWER: 12,5%